### v3

# Is Doublet-Triplet Splitting Necessary?

Abstract
We demonstrate that it is not necessary to substantially break the mass degeneracy between the Standard Model Higgs boson doublet and the corresponding scalar leptoquark, where these two fields comprise a single five-dimensional $SU(5)$ representation. More precisely, we show that the experimental data on partial proton decay lifetimes cannot place any meaningful bound on the mass of the scalar leptoquark, a.k.a. the color triplet, if one includes effects of higher-dimensional operators within the Georgi-Glashow model. We also point out that the complete suppression of proton decay via the scalar leptoquark mediation is compatible with the partial suppression of the gauge mediated two-body proton decay signatures. Finally, we derive a proper limit on the cutoff scale of the aforementioned higher-dimensional operators and outline how this scenario could be tested.
I Introduction
One of the most prominent features of the the Georgi-Glashow model Georgi:1974sy is an ever-present need to introduce an enormous splitting between the masses of the two multiplets residing in the same five-dimensional scalar representation $5_{H}$ of the theory. One of these two multiplets is the Standard Model (SM) Higgs boson doublet $D$, whereas the other one is the color triplet $T$ that generates proton decay. Since the Higgs boson is light while the proton decay mediating scalar leptoquark should be extremely heavy, one needs to strongly break mass degeneracy between them. This issue has been referred to in the literature as the ($D$)oublet-($T$)riplet splitting problem.
We demonstrate that it is not necessary to accomplish the aforementioned mass split if one includes effects of the higher-dimensional operators within the Georgi-Glashow model. The inclusion of these operators allows one to completely suppress proton decay signatures via the scalar leptoquark mediation without affecting viability of the mechanism that generates masses of the SM charged fermions. We furthermore show that this intriguing possibility is also compatible with the partial suppression of the gauge mediated two-body proton decay signatures.
We introduce the most relevant details of the $SU(5)$ setup and the accompanying notation in Sec. II. The proton decay suppression mechanisms for both the scalar and gauge mediations are discussed in detail in Sec. III. We subsequently show how to establish an accurate upper bound on the cutoff scale associated with the higher-dimensional operators if one resorts to these suppression mechanisms in Sec. IV. Final remarks are presented in Sec. V and all our findings summarized in Sec. VI.
II Setup
We consider the following $d=4$ and $d=5$ operators
$\mathcal{L}_{Y}=10_{F}^{\alpha ij}\bigg{\{}Y_{d}^{\alpha\beta}\overline{5}_{Fi}^{\beta}5^{*}_{Hj}+\frac{1}{\Lambda}Y_{1}^{\alpha\beta}\overline{5}_{Fi}^{\beta}5^{*}_{Hk}24_{Hj}^{k}$
$+\frac{1}{\Lambda}Y_{2}^{\alpha\beta}\overline{5}_{Fk}^{\beta}5^{*}_{Hi}24_{Hj}^{k}\bigg{\}}+10_{F}^{\alpha ij}10_{F}^{\beta kl}5_{H}^{m}\bigg{\{}Y_{u}^{\alpha\beta}\epsilon_{ijklm}$
$+\frac{1}{\Lambda}Y_{3}^{\alpha\beta}24_{Hm}^{n}\epsilon_{ijkln}+\frac{1}{\Lambda}Y_{4}^{\alpha\beta}24_{Hi}^{n}\epsilon_{jklmn}\bigg{\}}+\mathrm{h.c.},$ (1)
where $\Lambda$ is a cutoff scale of the theory, $\alpha,\beta=1,2,3$ are family indices, $i,j,k,l,m,n=1,\ldots,5$ are $SU(5)$ indices, and $Y_{d}$, $Y_{1}$, $Y_{2}$, $Y_{u}$, $Y_{3}$, and $Y_{4}$ are Yukawa coupling matrices. Here, scalars comprise $5_{H}$ and $24_{H}$, the SM fermions reside in $\overline{5}_{F}^{\alpha}$ and $10_{F}^{\alpha}$, while gauge fields are accommodated in $24_{G}$ Georgi:1974sy .
The first bracket of Eq. (1) contains operators that generate masses of the charged leptons and down-type quarks, whereas the operators from the second bracket of Eq. (1) affect exclusively the up-type quark masses.
The vacuum expectation values (VEVs) of relevance are
$\langle 24_{H}\rangle$ $=v_{24}\mathrm{diag}\left(-1,-1,-1,3/2,3/2\right),$ (2)
$\langle 5_{H}\rangle$ $=(0\quad 0\quad 0\quad 0\quad v_{5}/\sqrt{2})^{T},$ (3)
where $v_{5}\approx 246$ GeV correctly reproduces masses of the SM gauge boson fields $W_{\mu}^{\pm 1}\in 24_{G}$ and $Z^{0}_{\mu}\in 24_{G}$. (The field superscripts denote electric charges in units of the positron charge.)
Unification of the SM gauge couplings at scale $M_{\mathrm{GUT}}$, if successful, stipulates that
$M_{\mathrm{GUT}}\equiv M_{X}=M_{Y}=\sqrt{25/8}g_{\mathrm{GUT}}v_{24},$ (4)
where $g_{\mathrm{GUT}}$ is a gauge coupling constant at $M_{\mathrm{GUT}}$, while $M_{X}$ and $M_{Y}$ are masses of the proton decay mediating gauge fields $X_{\mu}^{\pm 4/3}\in 24_{G}$ and $Y_{\mu}^{\pm 1/3}\in 24_{G}$, respectively.
$\mathcal{L}_{Y}$ yields the following charged fermion mass matrices
$M_{E}$ $=v_{5}\bigg{\{}\frac{1}{2}Y_{d}+\frac{3}{4}Y_{1}\epsilon-\frac{3}{4}Y_{2}\epsilon\bigg{\}},$ (5)
$M_{D}$ $=v_{5}\bigg{\{}\frac{1}{2}Y_{d}^{T}+\frac{3}{4}Y_{1}^{T}\epsilon+\frac{1}{2}Y_{2}^{T}\epsilon\bigg{\}},$ (6)
$M_{U}$ $=v_{5}\bigg{\{}\sqrt{2}\left(Y_{u}+Y_{u}^{T}\right)+\frac{3}{\sqrt{2}}\left(Y_{3}+Y_{3}^{T}\right)\epsilon$
$+\left(\frac{1}{2\sqrt{2}}Y_{4}-\sqrt{2}Y_{4}^{T}\right)\epsilon\bigg{\}},$ (7)
where we introduce a dimensionless parameter $\epsilon=v_{24}/\Lambda$.
Mass matrices of the SM fermions are diagonalized via
$E^{T}_{c}M_{E}E=M_{E}^{\mathrm{diag}},$ (8)
$D^{T}_{c}M_{D}D=M_{D}^{\mathrm{diag}},$ (9)
$U^{T}_{c}M_{U}U=M_{U}^{\mathrm{diag}},$ (10)
$N^{T}M_{N}N=M_{N}^{\mathrm{diag}},$ (11)
where $E_{c}$, $E$, $D_{c}$, $D$, $U_{c}$, $U$, and $N$ are $3\times 3$ unitary matrices. $M_{N}$ is a $3\times 3$ mass matrix for neutrinos, where we explicitly assume neutrinos to be Majorana particles.
III Proton Decay Suppression
The only scalar that mediates proton decay in this scenario is $T^{-1/3}_{i}\equiv T_{i}\in 5_{H}^{i}$, where $i=1,2,3$. Its interactions with the SM fermions, as derived from Eq. (1), are
$(i)\;u^{T}_{k,\alpha}C^{-1}e_{\beta}T^{*}_{k}:$
$\bigg{\{}U^{T}\bigg{[}-\frac{Y_{d}}{\sqrt{2}}+\frac{Y_{1}}{\sqrt{2}}\epsilon+\frac{3}{2\sqrt{2}}Y_{2}\epsilon\bigg{]}E\bigg{\}}_{\alpha\beta},$ (12)
$(ii)\;\epsilon_{ijk}u^{C,T}_{i,\alpha}C^{-1}d^{C}_{j,\beta}T^{*}_{k}:$
$\bigg{\{}U_{c}^{\dagger}\bigg{[}\frac{Y_{d}}{\sqrt{2}}-\frac{Y_{1}}{\sqrt{2}}\epsilon+\frac{Y_{2}}{\sqrt{2}}\epsilon\bigg{]}D_{c}^{*}\bigg{\}}_{\alpha\beta},$ (13)
$(iii)\;d^{T}_{k,\alpha}C^{-1}\nu_{\beta}T^{*}_{k}:$
$\bigg{\{}D^{T}\bigg{[}\frac{Y_{d}}{\sqrt{2}}-\frac{Y_{1}}{\sqrt{2}}\epsilon-\frac{3}{2\sqrt{2}}Y_{2}\epsilon\bigg{]}N\bigg{\}}_{\alpha\beta},$ (14)
$(iv)\;\epsilon_{ijk}u^{T}_{i,\alpha}C^{-1}d_{j,\beta}T_{k}:$
[…]
$(v)\;u^{C,T}_{k,\alpha}C^{-1}e^{C}_{\beta}T_{k}:$
[…]
Since the linear combinations of Yukawa coupling matrices that enter the charged fermion masses $M_{E}$, $M_{D}$, and $M_{U}$ differ from those that are featured in the interactions of the scalar leptoquark $T$ with the SM fermions, one can suppress the latter without affecting viability of the former. For example, one can set to zero all quark-quark couplings of $T$, as given in Eqs. (13) and (15), by imposing the following three relations
$Y_{2}\epsilon$ $=Y_{1}\epsilon-Y_{d}$ (17)
$Y_{u}+Y_{u}^{T}$ $=(Y_{3}+Y_{3}^{T})\epsilon\equiv Y_{S},$ (18)
$Y_{4}$ $=-Y_{4}^{T}\equiv Y_{A},$ (19)
where the viability of Eq. (18) will be discussed in detail later on. These assumptions consequentially lead to
$Y_{d}$ $=\frac{4}{5v_{5}}E^{*}_{c}M_{E}^{\mathrm{diag}}E^{\dagger},$ (20)
$Y_{1}$ $=\frac{4}{5\epsilon v_{5}}D^{*}M_{D}^{\mathrm{diag}}D^{\dagger}_{c},$ (21)
$Y_{S}$ $=\frac{\sqrt{2}}{10v_{5}}\left(U^{*}_{c}M_{U}^{\mathrm{diag}}U^{\dagger}+U^{*}M_{U}^{\mathrm{diag}}U_{c}^{\dagger}\right),$ (22)
$Y_{A}$ $=\frac{\sqrt{2}}{5\epsilon v_{5}}\left(U^{*}_{c}M_{U}^{\mathrm{diag}}U^{\dagger}-U^{*}M_{U}^{\mathrm{diag}}U_{c}^{\dagger}\right).$ (23)
Even though we still need to demonstrate that one can indeed cancel $d=4$ contributions with $d=5$ contributions in the up-type quark sector with perturbative Yukawa couplings, Eqs. (17) through (23) already imply that generation of viable charged fermion masses is compatible with complete suppression of scalar mediated proton decay signatures. Therefore, one of the main results of this manuscript is that experimental data on partial proton decay lifetimes cannot provide any meaningful constraint on the scalar leptoquark mass if one includes effects of higher-dimensional operators. Our result is applicable to any non-supersymmetric $SU(5)$ setup that ($i$) uses a single $5_{H}$ and ($ii$) resorts to $d=5$ operators to provide viable masses for the SM charged fermions.
The next question we want to address is whether the complete suppression of proton decay signatures via the scalar leptoquark mediation is compatible with the partial suppression of the gauge mediated proton decay.
To that end, we recall that the relevant $d=6$ operators, which govern all eight two-body proton decay channels via exchanges of $X_{\mu}^{\pm 4/3}$ and $Y_{\mu}^{\pm 1/3}$ gauge bosons are FileviezPerez:2004hn ; Nath:2006ut
$\mathcal{O}_{I}=\frac{g_{\mathrm{GUT}}^{2}}{2M_{\mathrm{GUT}}^{2}}c(e^{C}_{\alpha},d_{\beta})\epsilon_{ijk}\overline{u^{C}_{i}}\gamma^{\mu}u_{j}\overline{e^{C}_{\alpha}}\gamma_{\mu}d_{k\beta},$ (24)
$\mathcal{O}_{II}=\frac{g_{\mathrm{GUT}}^{2}}{2M_{\mathrm{GUT}}^{2}}c(e_{\alpha},d^{C}_{\beta})\epsilon_{ijk}\overline{u^{C}_{i}}\gamma^{\mu}u_{j}\overline{d^{C}_{k\beta}}\gamma_{\mu}e_{\alpha},$ (25)
$\mathcal{O}_{III}=\frac{g_{\mathrm{GUT}}^{2}}{2M_{\mathrm{GUT}}^{2}}c(\nu_{\rho},d_{\alpha},d^{C}_{\beta})\epsilon_{ijk}\overline{u^{C}_{i}}\gamma^{\mu}d_{j\alpha}\overline{d^{C}_{k\beta}}\gamma_{\mu}\nu_{\rho}.$ (26)
The dimensionless coefficients of interest read
$c(e^{C}_{\alpha},d_{\beta})=\left(U_{c}^{\dagger}U\right)_{11}\left(E_{c}^{\dagger}D\right)_{\alpha\beta}+(U_{c}^{\dagger}D)_{1\beta}(E_{c}^{\dagger}U)_{\alpha 1},$ (27)
$c(e_{\alpha},d_{\beta}^{C})=\left(U_{c}^{\dagger}U\right)_{11}\left(D_{c}^{\dagger}E\right)_{\beta\alpha},$ (28)
$c(\nu_{\rho},d_{\alpha},d^{C}_{\beta})=\left(U_{c}^{\dagger}D\right)_{1\alpha}(D_{c}^{\dagger}N)_{\beta\rho}.$ (29)
Although proton decay via gauge boson exchange cannot be completely rotated away in $SU(5)$ Nandi:1982ew , it is possible to significantly suppress coefficients in Eqs. (27), (28), and (29Dorsner:2004xa . This, for example, can be achieved with the followings set of conditions Dorsner:2004xa
$\left(U^{\dagger}_{c}D\right)_{1\alpha}=0,$ (30)
$\left(E^{\dagger}_{c}D\right)_{1\alpha}=\left(E^{\dagger}_{c}D\right)_{\alpha 1}=0,$ (31)
$\left(D^{\dagger}_{c}E\right)_{1\alpha}=\left(D^{\dagger}_{c}E\right)_{\alpha 1}=0,$ (32)
where $\alpha=1,2$. Namely, since Eq. (30) implies, via $U^{\dagger}D=\mathrm{diag}(e^{i\phi_{1}},e^{i\phi_{2}},e^{i\phi_{3}})V_{\mathrm{CKM}}\mathrm{diag}(e^{i\phi_{4}},e^{i\phi_{5}},1)$, that $|(U_{c}^{\dagger}U)_{11}|=|(V_{\mathrm{CKM}})_{13}|$, one finds that Eqs. (30), (31), and (32) facilitate the following levels of suppression
$\max|c(e^{C}_{\alpha},d_{\beta})|=2\,\,\rightarrow\,\,\max|c(e^{C}_{\alpha},d_{\beta})|=|(V_{\mathrm{CKM}})_{13}|,$
$\max|c(e_{\alpha},d_{\beta}^{C})|=1\,\,\rightarrow\,\,\max|c(e_{\alpha},d_{\beta}^{C})|=|(V_{\mathrm{CKM}})_{13}|,$
$\max|c(\nu_{\rho},d_{\alpha},d^{C}_{\beta})|=1\,\,\rightarrow\,\,\max|c(\nu_{\rho},d_{\alpha},d^{C}_{\beta})|=0.$
Here, $V_{\mathrm{CKM}}$ represents the Cabibbo-Kobayashi-Maskawa mixing matrix with $|(V_{\mathrm{CKM}})_{13}|\approx 4\times 10^{-3}$, while $\phi_{1}$, $\phi_{2}$, $\phi_{3}$, $\phi_{4}$, and $\phi_{5}$ are arbitrary phases.
Even though the matrix elements of $Y_{d}$, $Y_{1}$, $Y_{S}$, and $Y_{A}$, as given in Eqs. (20), (21), (22), and (23), respectively, become somewhat more constrained after implementation of Eqs. (30), (31), and (32), it is still possible to completely suppress scalar mediated proton decay signatures. Again, Eq. (30) establishes relation between unitary matrices $U_{c}$ and $U$ via $V_{\mathrm{CKM}}$, whereas Eqs. (31) and (32) connect unitary transformations in the charged lepton and down-type quark sectors. For example, Eq. (31) relates $E_{c}$ and $D$ via
$E_{c}=D\begin{pmatrix}0&0&e^{i\xi_{1}}\\ 0&e^{i\xi_{2}}&0\\ e^{i\xi_{3}}&0&0\end{pmatrix},$ (33)
where $\xi_{1}$, $\xi_{2}$, and $\xi_{3}$ are arbitrary phases. Eq. (32) introduces an analogous relationship between $D_{c}$ and $E$.
IV Cutoff Scale Limit
Let us now derive an accurate upper bound on the cutoff scale $\Lambda$ of the setup that addresses charged fermion masses via Eq. (1) if one furthermore implements Eqs. (30), (31), and (32) to partially suppress gauge mediated proton decay signatures. This result improves the findings of Ref. Dorsner:2006hw , where the need to use $d=5$ operators to generate experimentally observed mismatch between the masses of charged leptons and down-type quarks and conditions of Eqs. (31) and (32) were exploited to find the associated upper bound on the cutoff scale. We, on the other hand, find that the most stringent limit on $\Lambda$ originates from implementation of Eq. (30) and the fact that the leading term in the up-type quark mass matrix $M_{U}$ is symmetric in the flavor space, as evident from Eq. (7).
Since a symmetric form of the up-type quark mass matrix would imply $|(U_{c}^{\dagger}U)_{11}|=1$, whereas Eq. (30) yields $|(U_{c}^{\dagger}U)_{11}|=|(V_{\mathrm{CKM}})_{13}|$, one needs to have a substantial skew-symmetric component in $M_{U}$. To quantify this, we decompose $M_{U}$ into a sum of a symmetric part $S$ and a skew-symmetric part $A$ via
$M_{U}=U^{*}_{c}M_{U}^{\mathrm{diag}}U^{\dagger}=\frac{v_{5}}{\sqrt{2}}\left(S+A\right).$ (34)
For our purposes it is sufficient to take $M_{U}^{\mathrm{diag}}=\mathrm{diag}(0,0,m_{t})=\mathrm{diag}(0,0,y_{t}v_{5}/\sqrt{2})$, where $m_{t}$ and $y_{t}$ are the top quark mass and Yukawa coupling at $M_{\mathrm{GUT}}$ scale, respectively. This decomposition allows us to find a lower limit on the largest possible entry in $A$, i.e., $\max|(A)_{\alpha\beta}|$, in units of $y_{t}$. Since the elements of $A$ can only originate from the very last operator in Eq. (7), our approach should yield an accurate lower limit on parameter $\epsilon\equiv v_{24}/\Lambda$ if we insist on perturbativity of Yukawa couplings. Coincidentally, decomposition of Eq. (34) is compatible with our assumptions of Eqs. (18) and (19). This, in turn, will enable us to explicitly demonstrate that one can cancel $d=4$ contributions with $d=5$ contributions with perturbative couplings in the up-type quark sector.
To perform our numerical analysis we first note that Eq. (30) implies that
$U_{c}=U\mathrm{diag}(e^{i\phi_{1}},e^{i\phi_{2}},e^{i\phi_{3}})V_{\mathrm{CKM}}\mathrm{diag}(e^{i\phi_{4}},e^{i\phi_{5}},1)X^{\dagger},$
where
$X=\begin{pmatrix}0&0&e^{i\eta_{4}}\\ e^{i(\eta_{1}/2+\eta_{2}+\eta_{3})}c_{\theta}&e^{i(\eta_{1}/2+\eta_{2}-\eta_{3})}s_{\theta}&0\\ -e^{i(\eta_{1}/2-\eta_{2}+\eta_{3})}s_{\theta}&e^{i(\eta_{1}/2-\eta_{2}-\eta_{3})}c_{\theta}&0\end{pmatrix}.$ (35)
Here, $c_{\theta}\equiv\cos\theta$ and $s_{\theta}\equiv\sin\theta$, where $\theta$ is a free parameter, while $\eta_{1}$, $\eta_{2}$, $\eta_{3}$, and $\eta_{4}$ are four arbitrary phases. We can, furthermore, represent unitary matrix $U$ via
$U=\mathrm{diag}(e^{i\beta_{1}},e^{i\beta_{2}},e^{i\beta_{3}})U_{23}U_{13}U_{12}\mathrm{diag}(e^{i\beta_{4}},e^{i\beta_{5}},1),$
where
$U_{12}=\begin{pmatrix}c_{u_{12}}&s_{u_{12}}&0\\ -s_{u_{12}}&c_{u_{12}}&0\\ 0&0&1\end{pmatrix},$ (36)
$U_{13}=\begin{pmatrix}c_{u_{13}}&0&s_{u_{13}}e^{-i\beta_{6}}\\ 0&1&0\\ -s_{u_{13}}e^{i\beta_{6}}&0&c_{u_{13}}\\ \end{pmatrix},$ (37)
$U_{23}=\begin{pmatrix}1&0&0\\ 0&c_{u_{23}}&s_{u_{23}}\\ 0&-s_{u_{23}}&c_{u_{23}}\\ \end{pmatrix}.$ (38)
To evaluate $U^{*}_{c}M_{U}^{\mathrm{diag}}U^{\dagger}$ of Eq. (34) and thus determine a lower limit on the largest entries in both $A$ and $S$, one needs to vary angle $\theta$ in $X$ and angles $u_{12}$, $u_{13}$, and $u_{23}$ in $U$. It also seems that one would need to vary phases $\eta_{1},\ldots,\eta_{4}$, $\phi_{1},\ldots,\phi_{5}$, and $\beta_{1},\ldots,\beta_{6}$. However, it turns out that $\eta_{4}$ drops out due to the form of $M_{U}^{\mathrm{diag}}$, while $\eta_{1}$, $\eta_{2}$, and $\eta_{3}$ can be absorbed through redefinition of phases $\phi_{4}$ and $\phi_{5}$. Also, any changes in $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ equally affect elements $(S+A)_{\alpha\beta}$ and $(S+A)_{\beta\alpha}$. This makes $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ completely irrelevant for our analysis. Finally, phases $\beta_{4}$ and $\beta_{5}$ that appear on the right-hand side of $M_{U}^{\mathrm{diag}}$ in Eq. (34) drop out from considerations, while $\beta_{4}$ and $\beta_{5}$ that appear on the left-hand side of $M_{U}^{\mathrm{diag}}$ can be absorbed in $\phi_{1}$ and $\phi_{2}$, respectively. After all is said and done, one only needs to vary four angles and five phases. These phases are $\phi_{1}$, $\phi_{2}$, $\phi_{4}$, $\phi_{5}$, and $\beta_{6}$, where $\phi_{3}$ is also irrelevant as it can be interpreted as an overall phase in Eq. (34), upon redefinition of phases $\phi_{1}$ and $\phi_{2}$.
We have accordingly performed a scan over $10^{6}$ configurations of the randomly chosen value sets of these nine parameters, i.e., $\theta$, $u_{12}$, $u_{13}$, $u_{23}$, $\phi_{1}$, $\phi_{2}$, $\phi_{4}$, $\phi_{5}$, and $\beta_{6}$, to find that $\max|(A)_{\alpha\beta}|\in(0.288y_{t},0.500y_{t})$ and $\max|(S)_{\alpha\beta}|\in(0.261y_{t},0.521y_{t})$. Moreover, $\max|(A)_{\alpha\beta}|$ and $\max|(S)_{\alpha\beta}|$ tend to be correlated. Namely, if the largest entry in $A$ is on the larger side, so is the largest entry of matrix $S$. To that end we present in Fig. 1 a plot of $\max|(A)_{\alpha\beta}|$ vs. $\max|(S)_{\alpha\beta}|$ for $10^{4}$ randomly chosen sets of values of four angles and five phases of relevance.
Figure 1: A plot of $\max|(A)_{\alpha\beta}|$ vs. $\max|(S)_{\alpha\beta}|$ for $10^{4}$ sets of randomly chosen values of relevant parameters. For demonstration purpose we set $y_{t}=1$.
To find an upper bound on $\Lambda$ we first observe that the relevant limits are $\max|(A)_{\alpha\beta}|>0.288y_{t}$ and $\max|(S)_{\alpha\beta}|>0.261y_{t}$. Also, the assumption that $Y_{4}$ is skew-symmetric, is consistent with the search for an upper limit on $\Lambda$ as it yields the smallest possible maximum for values of elements in $Y_{4}$ matrix. We can thus write, if we take $Y_{4}=-Y_{4}^{T}$ and use Eqs. (7) and (34) to find both the symmetric and skew-symmetric contributions in $M_{U}$, that $S=2(Y_{u}+Y_{u}^{T})+3(Y_{3}+Y_{3}^{T})\epsilon$ and $A\equiv 5/2Y_{4}\epsilon$. The latter identity, then, leads to the following inequality
$\frac{5}{2}\max|(Y_{4})_{\alpha\beta}|\epsilon\equiv\frac{5}{2}\max|(Y_{4})_{\alpha\beta}|\frac{v_{24}}{\Lambda}>0.288y_{t}.$ (39)
To proceed, one would need to establish what the viable values of $v_{24}$ and $y_{t}$, to be used as input in Eq. (39), are. We accordingly observe that $v_{24}=M_{\mathrm{GUT}}\sqrt{(2\alpha_{\mathrm{GUT}}^{-1})/25\pi}$, where $\alpha_{\mathrm{GUT}}=g_{\mathrm{GUT}}^{2}/(4\pi)$. Also, a model Dorsner:2005fq ; Dorsner:2005ii ; Dorsner:2006hw that $(i)$ uses perturbative couplings to generate neutrino masses, $(ii)$ relies on operators of Eq. (1) to generate charged fermion masses, and $(iii)$ implements Eqs. (30), (31), and (32) to suppress gauge mediated proton decay signatures, yields $\alpha_{\mathrm{GUT}}^{-1}=37.1$ and $y_{t}=0.42$, where $\alpha_{\mathrm{GUT}}$ is evaluated via the two-loop level gauge coupling unification analysis, whereas $y_{t}$, at $M_{\mathrm{GUT}}$, is obtained through the one-loop level renormalisation group equation running. If these values for $\alpha_{\mathrm{GUT}}$ and $y_{t}$ are inserted into Eq. (39) we find that the upper limit on $\Lambda$ for a model of Ref. Dorsner:2005fq reads $\Lambda<57M_{\mathrm{GUT}}$, where we assume that $5/2\max|(Y_{4})_{\alpha\beta}|=\sqrt{4\pi}$. This bound on $\Lambda$, as inferred from Eq. (39), is another key finding of this manuscript. It is, as already advocated, much more stringent than the $\Lambda\lesssim 900M_{\mathrm{GUT}}$ bound reported in Ref. Dorsner:2006hw .
It is now trivial to show that it is possible to cancel $d=4$ contributions with $d=5$ terms in the up-type quark sector, as required by Eq. (18), if one is to completely suppress scalar mediated proton decay signatures. Namely, if we insert conditions of Eqs. (18) and (19) into Eq. (7) and implement numerical analysis of the decomposition of Eq. (34), we find that
$5\max|(Y_{3}+Y_{3}^{T})_{\alpha\beta}|\epsilon>0.261y_{t}.$ (40)
Since the right-hand side of Eq. (39) is larger than the one in Eq. (40), we see that one can simultaneously accomplish complete suppression of the scalar mediated proton decay and partial suppression of the gauge mediated proton decay, where the correct upper bound on the cutoff scale $\Lambda$ in Eq. (39) originates from Eqs. (7) and (30).
Let us, as a qualitative example, assume that $\epsilon=0.05$. This, for $\alpha_{\mathrm{GUT}}^{-1}=37.1$ and $y_{t}=0.42$, implies that $\Lambda=19.4M_{\mathrm{GUT}}$, $2.25\max|(Y_{3}+Y_{3}^{T})_{\alpha\beta}|>$ 2.2, $2.45\max|(Y_{4})_{\alpha\beta}|/2>$ 2.4, and $0.882\max|(Y_{u}+Y_{u}^{T})_{\alpha\beta}|>$ 0.88 if one is to implement suppression of both scalar and gauge mediations of proton decay.
If one only wants to suppress scalar mediated proton decay, one can completely drop a skew-symmetric contribution towards $M_{U}$ in Eq. (7). This could be accomplished by setting, for example, $Y_{4}=0$. In this scenario, $M_{U}=M_{U}^{T}$ implies $U_{c}\equiv U$. We can thus resort to the decomposition of Eq. (34), drop $A$, set $U_{c}\equiv U$, and vary three angles in $U$, together with three phases $\beta_{4}$, $\beta_{5}$, and $\beta_{6}$, to establish correct upper limit on the cutoff scale $\Lambda$. This procedure, combined with the condition of Eq. (18), yields
$5\max|(Y_{3}+Y_{3}^{T})_{\alpha\beta}|\epsilon>0.334y_{t}.$ (41)
It is clear that if one drops a skew-symmetric contribution to $M_{U}$, one would obtain somewhat more stringent limit on the cutoff scale $\Lambda$. It is thus preferable to have the $d=5$ skew-symmetric contribution towards $M_{U}$ or even obligatory if one is to suppress proton decay signatures through gauge boson mediation. Amazingly, though, it is entirely possible to completely suppress proton decay through scalar mediation even when the up-type quark mass matrix is symmetric as long as $\Lambda$ satisfies Eq. (41).
V Final Remarks
Final comments are in order.
  • Our derivation of the limit on $\Lambda$ is accurate up to the corrections of the order of $m_{c}/m_{t}$, where $m_{c}$ is the charm quark mass. Moreover, the values of elements in $V_{\mathrm{CKM}}$ used to generate bounds in Eqs. (39), (40), and (41) are sourced from the Particle Data Group ParticleDataGroup:2022pth , although the only $V_{\mathrm{CKM}}$ entry that matters for our analysis is the Cabibbo angle.
  • We have explicitly demonstrated how to suppress the quark-quark couplings of the scalar leptoquark $T$ in Sec. III. We could have alternatively opted to suppress the quark-lepton couplings of Eqs. (12), (14), and (16) via $Y_{2}\epsilon=\frac{2}{3}\left(Y_{d}-Y_{1}\epsilon\right)$ and $\left(Y_{3}+Y_{3}^{T}\right)\epsilon=\left(Y_{u}+Y_{u}^{T}\right)+\left(\frac{3}{2}Y_{4}-Y_{4}^{T}\right)\epsilon$ in order to eliminate proton decay through the color triplet mediation. Note, however, that we cannot simultaneously suppress both sets of the scalar leptoquark couplings. This distinguishes our proposal from an approach advocated in Ref. Dvali:1992hc that insists on complete decoupling of the color triplet from the SM fermions, thus resulting, as specified in Ref. Dvali:1995hp , in existence of long-lived colored state(s).
  • If the color triplet $T\in 5_{H}$ is indeed light, it would form a complete $SU(5)$ multiplet with the SM Higgs boson doublet $D\in 5_{H}$ as far as the analysis of the running of the gauge coupling constants is concerned. Since the SM Higgs boson doublet is always beneficial for unification and aids the largeness of $M_{\mathrm{GUT}}$, the lightness of $T$ would consequentially lower maximal possible value of $M_{\mathrm{GUT}}$, thus making the proton decay signatures through the gauge boson mediation more accessible.
  • Suppression of the scalar mediation of proton decay signatures somewhat constrains Yukawa couplings of leptoquark $T$ to the SM fermions. (See, for example, Eqs. (20), (21), (22), and (23).) If one, furthermore, suppresses proton decay through gauge mediation, one would also introduce strong correlation between pairs of unitary matrices ($U$, $U_{c}$), ($D$, $E_{c}$), and ($E$, $D_{c}$). Consequently, the detection of the scalar leptoquark $T$ at current or future colliders would enable the verification of consistency between lepton-quark-leptoquark or quark-quark-leptoquark interactions and the observed final states.
  • Our approach to the ($D$)oublet-($T$)riplet splitting problem can be easily applied within the supersymmetric framework. For example, within the minimal supersymmetric $SU(5)$ model, augmented with higher-dimensional operators to address observed charged fermion masses, one of the two triplets would have couplings given in Eqs. (12), (13), and (14), whereas the other triplet would have couplings given in Eqs. (15) and (16). One could then suppress either the quark-quark or lepton-quark interactions of both of these triplets using the conditions already discussed in our manuscript. This, in turn, would completely suppress all $d=5$ and $d=6$ proton decay operators associated with these two triplets. We defer a full viability study of cancellation between $d=4$ and $d=5$ contributions towards charged fermion masses that is required if one is to implement our approach towards the ($D$)oublet-($T$)riplet splitting problem within the minimal supersymmetric $SU(5)$ model with higher-dimensional operators to future publication. (For several attempts to address the ($D$)oublet-($T$)riplet splitting problem within the minimal supersymmetric $SU(5)$ setting see Refs. Berezhiani:1998hg ; Bajc:2002bv ; Bajc:2002pg .)
VI Conclusions
We demonstrate that without additional assumptions about the flavor structure of the theory, the experimental data on partial proton decay lifetimes cannot meaningfully constrain the mass of the scalar leptoquark if one includes effects of higher-dimensional operators. Our result is applicable to any non-supersymmetric $SU(5)$ setup that $(i)$ uses a single $5_{H}$, where the aforementioned leptoquark resides, and $(ii)$ resorts to the use of $d=5$ operators to provide viable masses for the SM charged fermions. This undermines the underlying premise behind the need to implement the doublet-triplet mass splitting in this class of models.
We also show that the complete suppression of proton decay via the scalar leptoquark mediation is compatible with the partial suppression of the gauge mediated proton decay signatures.
Finally, we show how to derive an accurate upper bound on the cutoff scale of the $d=5$ operators in the fermion sector if one implements partial suppression of the two-body proton decay signatures due to the gauge boson exchanges, complete suppression of the scalar mediated proton decay, or both.
Acknowledgments
I.D. acknowledges the financial support from the Slovenian Research Agency (research core funding No. P1-0035).